Mir Status 19 October 1997

NASA 6 Status Date: Sunday, October 19, 1997// MD 75/24

Mir Station System/Subsystem Activities

The Progress vestibule was successfully depressurized today. The station was commanded to free drift at DMT 11:37 this morning, and the crew then removed the Progress clamps and closed the Kvant-1 hatch. The depress was started at 13:05, and by the end of the comm pass the pressure had dropped to 37 mm. The valve was closed and the station returned to solar inertial attitude control at 13:41.

In further preparations for the IVA, the crew disconnected the cables for the Omega angular rate sensor in Kristall, and ORT-1 was enabled as the primary sensor, with ORT-2 in hot standby mode. They also disconnected the cables for the Kvant-2 Elektron, which receives power from Kristall. They shut down the Kvant-1 Elektron without purging after the 13:00 comm pass.

Mir Payload Activities
FB:HSR:I:23 CHAPAT Experiment Computer Monitoring
MG:BT:I:62 CAPE/MIM Daily Status Checks
MG:FAC:I:39 BEM/BSTC-M & BTR Facility Maintenance

Power Down of OPM and CHAPAT

OTHER ACTIVITIES

Dave reported yesterday that he used a pH strip from the Water kit and measured the liquid in the BTR. He said it registered 5.5 which is the same as other condensate measured in the station.

Experiment Operations -- Physical Fitness Assessment: Dave reported that the HRM procedures are needlessly complicated and iterative.

Earth Obs: Dave reported that he has been receiving the list of available sites and has had time to accomplish some observations.

SMP/Medical

David reported that he is not fatigued prior to the IVA tomorrow, iIim at 100%, and ready to go.i

Environmental Parameters (previous 24 hour period):
Parameter Min. Max. Nominal Range
P (mmHg) BB 723 729 660 - 850
ppO2 (mmHg) 165 168 140 - 200
ppCO2 (mmHg) 3.8 4.2 up to 8.0
ppH20 (mmHg) 13.2 14.3 8 - 14
Humidity (%) 49 54 30 - 75
Temp (xC) BB 26.4 27.5 18 - 28

Communications Summary / A/G with Consultant Group
09:54 - 10:00 IPCG

Unpowered times; PFC; HRM watch Radiogram; Earth Obs
1133 - 1134 IPCG

IVA Powered Hardware Operations (CAPE/MIM [powerdown prior to hatch closure])

Preparations for Spektr IVA - October 20, 1997
All payload activities have been successfully completed.

  • submit to reddit


Loading



Humans to Mars Summit 2019
Military Space Situational Awareness April 1 - April 2, 2019, London, UK
MilSatCom USA conference will return to Washington from the 26th – 27th June 2019
Space Access Conference 2019
Mission To Methonē by Les Johnson - Baen Books




Join our mailing list




Monthly Archives

About this Entry

This page contains a single entry by Keith Cowing published on October 19, 1997 4:27 PM.

Mir Status Report 5 October 1997 was the previous entry in this blog.

Mir Status 20 October 1997 is the next entry in this blog.

Find recent content on the main index or look in the archives to find all content.